The golden valley baseball team played 28 games.They won 3 times the number of games that they lost.How many games did they win?How many games did they lose?
Answer by TimothyLamb(4379)

W + L = 28
W = 3L
---
W + L = 28
3L + L = 28
4L = 28
---
L = 7
W = 21
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php